OpenStudy (carlyleukhardt):

I will medal! (x + 1) - (-2x - 5) A) 6 –x B) -x - 4 C) 3x - 4 D) 3x + 6

OpenStudy (johnn.):

Let's simplify step-by-step. x+1−(−2x−5) Distribute the Negative Sign: =x+1+−1(−2x−5) =x+1+−1(−2x)+(−1)(−5) =x+1+2x+5 Combine Like Terms: =x+1+2x+5 =(x+2x)+(1+5) =3x+6 Answer: =3x+6

OpenStudy (carlyleukhardt):

@JohnN. Thanks so much! <3x